C. Find the moment at Point A. -4 F1-6.6 lb, 1,-1.35 ft, l,-0.95 ft, l,-0.4 ft, 04-165°, α,-120°
Answer #1

F- G.6lb 02 х. data ,11: 1.35ft = 고65。 t80- = 180-165-150 e-180-X2 = 180-120-508 Nouo 021